OpenStudy (anonymous):

Solve the absoulte value inequailty below, expressing your answer in interval notation.

OpenStudy (anonymous):

\[\frac{ 1 }{ \left| x+5 \right| }\ge1\]

OpenStudy (anonymous):

I've gotten as far as \[-4\ge x\]

OpenStudy (anonymous):

and \[-6\le x\]

OpenStudy (anonymous):

the problem is that I can't figure out how to write my answer in interval notation

OpenStudy (anonymous):

\[-6 \le x \le -4\]

OpenStudy (anonymous):

I've gotten it that far, but interval notation is (?,?) or [?,?] or something similar.

OpenStudy (anonymous):

oh right sorry

OpenStudy (anonymous):

square brackets if it's less than/ greater than and equal to. round brackets if it's not equal to.

OpenStudy (anonymous):

< and > are round brackets, \[\le and \ge \]

OpenStudy (anonymous):

* are square brackets

OpenStudy (anonymous):

So what should the answer be?

OpenStudy (anonymous):

[-6,-4] is what I've come up with, but my online program says that is incorrect.

OpenStudy (anonymous):

Well that is in interval notation. are you sure -6, -4 is correct?

OpenStudy (anonymous):

Yes

OpenStudy (rational):

Notice that the expression becomes undefined when \(x=-5\), so you need to exclude it from solution

OpenStudy (rational):

the solution after excluding would be \[\large [-6,-5) \cup (-5, -4]\]

OpenStudy (anonymous):

Thank you! I had tried a similar answer earlier but I must have used the wrong type of brackets.